OpenStudy (anonymous):

Find the equation of the line through the point (2, 5) that cuts off the least area from the first quadrant. Give your answer using the form below. y − A = B(x − C)

OpenStudy (anonymous):

lets say the equation is y = mx+b m is the slope, b is the y intercept. the line will cut a triangular region in the first quadrant. its area is 1/2 base times height. the height is the y intercept and the base is y intercept divided by slope so the area is b^2/2m

OpenStudy (anonymous):

now, we know that the line passes through (2,5) so 5 = 2m+b, or 2m = 5-b substitute that in the area (call it A) A = b^2/5-b to find the least area, differentiate the area with respect to the height and equate it to 0 dA/db = 0 find b use that to find m you have the equation of the line manipulate that to get it in the form that the problem asks for.

OpenStudy (anonymous):

got it?